How do You solve the system below by elimination?
6x+4y= 132
x+y= 25

Answer by ankor@dixie-net.com(22740)

How do You solve the system below by elimination?
6x+4y= 132
x+y= 25
Multiply the 2nd equation by -4, you have:
6x + 4y = 132
-4x - 4y = -100
-------------------addition eliminates y, find x
2x = 32
x = 16
I'll let you find y